oliviac1111:

Arturo is building a flower bed in the shape of a right triangle. The hypotenuse of the right triangle is 13 feet. One of the legs needs to be 7 feet longer than the other leg. Which equation can be used to find x, the length of the shorter leg? x(x + 7) = 13 x(x + 7) = 169 x2 + (x + 7)2 = 13 x2 + (x + 7)2 = 169

xxemilianaxx:

Well what do you think?

supie:

Asked & Answered here.

oliviac1111:

@supie wrote:
Asked & Answered here.
thank you i failed the last assignment i didnt realize it was the same question

oliviac1111:

im not sure

Florisalreadytaken:

 \[\color{red}c= 13 \ \ \ \ ; \ \ \ \ \color{orange} a =x \ \ \ ; \ \ \ \ \color{cyan}b=(x+7) \] \[\color{red}{c^2} = \color{orange}{a^2} + \color{cyan}{b^2}\] can you now?

oliviac1111:

wait so this is pythagorean theorem

Florisalreadytaken:

yes!

oliviac1111:

so i would do 13=x+7+x? im really good at simple problems in Pythagorean theorem but not this

Florisalreadytaken:

no, youre a bit off it would be \[ \color{red}{13}^2 = \color{orange}{x}^2 + (\color{cyan}{x+7})^2 ​\] we can also write that as \( \color{orange}{x}^2 + (\color{cyan}{x+7})^2 = \color{red}{13}^2 \) right? which option is that (so not forget to do the \(13^2\))

oliviac1111:

wait huh wich option is what

Florisalreadytaken:

which option is: \[ \color{orange}{x}^2 + (\color{cyan}{x+7})^2 = \color{red}{169} \]

oliviac1111:

b i think

Florisalreadytaken:

how about you have a look at the last one?

oliviac1111:

so its d